### v3

# Renato Umeton1,4, Amelia Fiske 2, Leo Anthony Celi1,5

Abstract
Large-scale clinical databases offer opportunities for medical research, but their complexity creates barriers to effective use. The Medical Information Mart for Intensive Care (MIMIC-IV), one of the world’s largest open-source electronic health record databases, traditionally requires both SQL proficiency and clinical domain expertise. We introduce M3, a system that enables natural language querying of MIMIC-IV data through the Model Context Protocol. With a single command, M3 retrieves MIMIC-IV from PhysioNet, launches a local SQLite instance or connects to hosted BigQuery, and allows researchers to pose clinical questions in plain English. We evaluated M3 using one hundred questions from the EHRSQL 2024 benchmark with two language models: the proprietary Claude Sonnet 4 achieved 94% accuracy, while the open-source gpt-oss-20B (deployable locally on consumer hardware) achieved 93% accuracy. Both models translate natural language into SQL, execute queries against MIMIC-IV, and return structured results alongside the underlying query for verification. Error analysis revealed that most failures stemmed from complex temporal reasoning or ambiguous question phrasing rather than fundamental architectural limitations. The comparable performance of a smaller open-source model demonstrates that privacy-preserving local deployment is viable for sensitive clinical data analysis. M3 lowers technical barriers to critical care data analysis while maintaining security through OAuth2 authentication, query validation, and comprehensive audit logging.
1 Introduction
1.1 The Challenge of Analyzing Large-Scale Clinical Databases
The digital transformation of healthcare has led to the generation and accumulation of vast quantities of electronic health record (EHR) data [evans2016electronic], creating invaluable resources for secondary use of data, such as medical research that offer deep insights into disease patterns, treatment efficacy, and patient outcomes. However, the barrier to use these datasets is often high, due to data inherent complexity and required data querying skills. More in detail, clinical databases are typically relational, consisting of numerous interconnected tables and a multitude of fields, each with specific definitions and coding schemes. Effectively navigating and extracting meaningful information from such intricate structures necessitates specialized technical skills, primarily proficiency in SQL, a thorough understanding of the database schema, and how data points are temporarily connected and semantically linked. The SQL technical requirement forms a substantial barrier for many clinical researchers, while the clinical data domain poses a barrier to entry to most non-clinical data scientists or analysts. Consequently, the technical skill set needed to directly query complex databases like MIMIC-IV [johnson2023mimiciv] can limit the pool of researchers able to leverage these resources, potentially impeding the pace of innovation (e.g., clinical process improvement). This also highlights an interdisciplinary gap where clinical experts, who formulate the critical research questions, may be disconnected from the data extraction process, which often falls to data scientists or programmers. Tools that can bridge this divide by simplifying data access are therefore of growing importance, and some are already in use in academic medical centers and other clinical settings. [statnewsSvelte, jiang2023health, doi:10.1056/AIcs2300191]
Anthropic’s Model Context Protocol (MCP) [anthropic2024mcp] provides a standardized framework for managing AI model interactions with external software tools and data sources, offering a promising approach to address these accessibility challenges through secure and controlled interfaces.
1.2 The Role of MIMIC-IV in Critical Care Research
MIMIC-IV stands as a cornerstone of publicly available database for critical care research. Developed by the MIT Laboratory for Computational Physiology, this dataset contains de-identified health data associated with patients admitted to intensive care units (ICUs) or the emergency department at the Beth Israel Deaconess Medical Center. MIMIC-IV (version 3.1) [mimicivphysionet] includes data from approximately 364,627 unique individuals (each represented by a unique subject_id), 546,028 hospitalizations and 94,458 unique ICU stays. The dataset is rich in detail, including patient demographics, vital sign measurements, laboratory test results, medications, procedures, and more.
MIMIC-IV is widely utilized in the research community for developing and validating clinical prediction models, understanding disease trajectories, evaluating treatment interventions, and ultimately aiming to improve patient care in critical settings. The availability of MIMIC-IV through the PhysioNet platform [mimicivphysionet], which provides access modalities such as Google BigQuery for the full dataset, enhances research transparency and reproducibility, key elements of scientific progress [goldberger2000physionet]. The public, albeit credentialed, nature of MIMIC-IV enabled numerous research groups to work with standardized, high-fidelity clinical data, fostering collaboration and building upon prior work.
While large, the set of users would grow even larger, should MIMIC-IV data analysis carry a lower barrier to entry.
1.3 Introducing M3: Objectives and Contributions
This paper introduces M3, a project developed to address the challenges of accessing and analyzing MIMIC-IV data. The primary objective of M3 is to transform how researchers interact with this prime medical data resource by enabling natural language querying facilitated by AI assistance. Instead of writing complex SQL, users could pose questions in English and retrieve medical insights.
The key contributions of the M3 project are:
  • A novel software framework specifically designed to simplify data access for the MIMIC-IV database.
  • An architectural system, based on MCP, which facilitates interaction between AI agents and the MIMIC-IV data backend.
  • Demonstrated feasibility and performance through successful querying of both a small demo version of MIMIC-IV (using SQLite) and the full-scale dataset (using Google BigQuery).
  • A significant step towards lowering the technical barrier to entry for MIMIC-IV research, making the data more accessible to a broader range of researchers.
M3 represents a concrete application of Natural Language Interface (NLI) and text-to-SQL research, tailored to a specific, high-impact medical dataset, thereby moving from general research concepts to a practical, usable tool that significantly lowers the technical barrier to entry for MIMIC-IV research, making the data more accessible to a broader range of researchers while maintaining transparent and reproducible data provenance.
Figure 1: Results of a complex query, described in natural language as "Among patients who were diagnosed with anemia, unspecified since 2100, what are the top three most commonly prescribed medications that followed during the same hospital visit for patients in their 60 or above?"
We evaluate M3 using samples from the EHRSQL 2024 test set [ehrsql2024github], a benchmark designed for assessing natural language-to-SQL performance in clinical contexts. This dataset is based on the publicly available MIMIC-IV demo (v2.2) [johnson2023mimiciv_demo], which includes a representative subset of real hospital data and is freely accessible through PhysioNet. To illustrate the type of complex question that M3 can handle, Figure 1 shows the result of a query involving multiple temporal and clinical constraints. Specifically, this complex request inquires about the top three most commonly prescribed medications that were administered after a diagnosis of unspecified anemia (ICD code) during the same hospital admission, among patients aged 60 or older whose diagnosis occurred in the year 2100 or later. This result was obtained through the M3 system powered by Claude Sonnet 4  [anthropic2025claude4] via the MCP. The example highlights how M3 enables non-technical users to retrieve clinically meaningful insights from complex databases using only natural language.
For comparison, this is the corresponding correct query that the researcher should have entered otherwise:
Listing 1: Correct SQL query [ehrsql2024github]
SELECT T3.drug
FROM (
SELECT T2.drug, DENSE_RANK() OVER (ORDER BY COUNT(*) DESC) AS C1
FROM (
SELECT admissions.subject_id, diagnoses_icd.charttime, admissions.hadm_id
FROM diagnoses_icd
JOIN admissions ON diagnoses_icd.hadm_id = admissions.hadm_id
WHERE diagnoses_icd.icd_code = (
SELECT d_icd_diagnoses.icd_code
FROM d_icd_diagnoses
WHERE d_icd_diagnoses.long_title = ’anemia, unspecified’
AND strftime(’%Y’, diagnoses_icd.charttime) >= ’2100’
) AS T1
JOIN (
SELECT admissions.subject_id, prescriptions.drug, prescriptions.starttime, admissions.hadm_id
FROM prescriptions
JOIN admissions ON prescriptions.hadm_id = admissions.hadm_id
WHERE admissions.age >= 60
AND strftime(’%Y’, prescriptions.starttime) >= ’2100’
) AS T2
ON T1.subject_id = T2.subject_id
WHERE T1.charttime < T2.starttime
AND T1.hadm_id = T2.hadm_id
GROUP BY T2.drug
) AS T3
WHERE T3.C1 <= 3;
We also include a dedicated ethical considerations section  5 to reflect on the broader implications of lowering access barriers to clinical data via AI systems.
2 Related Work
2.1 Evolution of Clinical Database Access Tools
Recent years have seen significant progress in lowering the technical barriers to accessing and analyzing complex clinical databases, particularly for researchers without advanced programming expertise. Early efforts focused on direct SQL querying and basic graphical interfaces, requiring significant technical expertise from users. The MIMIC-II project [saeed2011mimicii] introduced web-based query builders and virtual machine environments, marking an important step toward simplifying database access for clinical researchers.
The development of MIMIC-IV expanded these capabilities through various access modalities, including cloud platforms such as Google BigQuery [johnson2022mimiciv, johnson2023mimiciv]. While this improved data accessibility and processing capabilities, the fundamental challenge of SQL expertise remained and was always compounded by the equally important required understanding of the clinical domain. Visual query builders and curated SQL templates [tian2024sqlucid] have attempted to bridge this gap, though often sacrificing query flexibility for ease of use.
The emergence of standards such as HL7 FHIR, the OMOP Common Data Model, and mCODE is enabling new, more scalable methods of accessing and sharing health data. The MIMIC-IV on FHIR implementation represents an important step toward standardized data access, though it brings its own complexities in terms of resource modeling and query patterns [hl7fhir, bennett2023mimic, ohdsiStandardization] .
2.2 Natural Language Interfaces for Medical Data
The development of natural language interfaces for databases (NLIDB) has seen several approaches evolve in parallel. Early NLIDB implementations on healthcare domain like MIMICSQL [wang2020text] demonstrated the basic feasibility of translating natural language to SQL, though they often struggled with query complexity and medical terminology variations. Subsequent systems such as EHRSQL [yu2023ehrsql] employed more sophisticated techniques to improve query understanding, showing better handling of medical terminology while still facing challenges with complex temporal relationships and nested queries common in clinical research.
2.3 Benchmarks and Evaluation Frameworks
The development of specialized benchmarks has been crucial for advancing the field. While general text-to-SQL benchmarks like BIRD [li2024can], Spider [yu2018spider] and WikiSQL [zhong2017seq2sql] provided foundational evaluation frameworks, they lack medical domain coverage and specificity. More recent efforts such as BiomedSQL [koretsky2025biomedsql] and the EHRSQL 2024 shared task [choi2024ehrsql] have introduced domain-specific challenges that better reflect real-world clinical querying needs. These benchmarks have revealed significant challenges in handling implicit medical knowledge, understanding temporal relationships in clinical data, managing hierarchical medical concepts, and integration with clinical workflows.
2.4 Security and Integration Frameworks
Security considerations in clinical database access have evolved from basic database-level security and input sanitization, as outlined in resources like the OWASP SQL Injection Prevention Cheat Sheet [owasp2023sqlcheat], to more comprehensive approaches. The introduction of the MCP  [anthropic2024mcp] represents a significant advance in AI-database integration that can support modern security standards, providing precise interaction patterns, access control mechanisms, audit capabilities, and reproducible query execution. Industry adoption of MCP has indeed grown across various domains including software development, scientific research, and biomedical [mcpharma2025], [mcpsuperagi2025].
2.5 Current Challenges and Opportunities
Existing solutions continue to face several key challenges. General-purpose text-to-SQL systems often struggle with medical terminology and relationships, while specialized medical systems may sacrifice query flexibility for security. Many current solutions lack robust mechanisms for ensuring query provenance and result reproduction. Technical integration requirements can remain substantial, and scaling to handle the complexity of full clinical databases presents ongoing challenges. To our knowledge, none of these is currently integrated in a desktop generative AI application, such as Claude Desktop [anthropic2025desktop] for instance.
M3 builds upon these foundations while addressing these challenges through its MCP-based architecture, specialized clinical tools, and robust security framework. By focusing specifically on the MIMIC-IV database and its unique characteristics, M3 aims to provide a more accessible yet secure approach to clinical data analysis.
3 Methodology
3.1 M3 Overview and System Architecture
M3 is designed as a robust, Python-based server application that facilitates natural language interaction with the MIMIC-IV critical care database. Its architecture (Figure 2) prioritizes secure, scalable, and user-friendly data access for clinical researchers.
The system employs a layered architecture comprising: (1) a data access layer supporting SQLite and BigQuery backends, (2) a security middleware implementing OAuth2 authentication and SQL validation, and (3) an MCP client built on the FastMCP framework that exposes tools to Large Language Model (LLM) agents. Standard software engineering best practices, such as (i) source code version control, (ii) modular architecture with abstract interfaces, (iii) functional and integration testing, are adopted across the project for ease of extension and support.
Figure 2: Conceptual Diagram of the M3 System Architecture
3.2 Data Sources and Access Layer
M3 supports two distinct backends for accessing the MIMIC-IV dataset, offering flexibility based on user needs and data scale:
  • Local SQLite Database: For rapid prototyping and development, M3 provides a local SQLite implementation using the official 100-patient demo subset of MIMIC-IV [johnson2023mimiciv]. This option requires minimal setup and incurs no cloud costs. The system handles the complete Extract, Transform, Load (ETL) process from PhysioNet data files to a local database, including schema inference and standardized null value handling.
  • Google BigQuery: For full-scale research, M3 connects to the complete MIMIC-IV v3.1 schemas [mimicivphysionet] on Google BigQuery. This implementation supports advanced features such as parameterized queries, cost estimation, and IAM-based access control. Access requires prior PhysioNet credentialing and an active Google Cloud project.
3.3 Configuration and Deployment
M3 provides an interactive shell interface for system configuration and management. Users can easily select their preferred backend (SQLite or BigQuery) and configure authentication settings through this interface. The system supports both interactive and programmatic configuration approaches, allowing flexible deployment options for different research environments.
3.4 Model Context Protocol (MCP) Tooling
M3 exposes its functionality through a two-tiered tool architecture compatible with the Model Context Protocol. These tools enable external LLM agents to translate natural language queries into structured database operations.
3.4.1 Core Database Tools
The foundational layer provides essential database access capabilities including schema discovery, table inspection, and query execution. These tools allow agents to understand the database structure and execute flexible and transparent SQL queries against the MIMIC-IV dataset.
3.4.2 Domain-Specific Clinical Tools
To reduce complexity for common clinical research patterns, M3 provides specialized tools that encapsulate frequent operations such as retrieving ICU stay information, laboratory results, and demographic distributions. These tools abstract complex joins and aggregations that would otherwise require extensive SQL expertise.
3.5 Security and Safeguards
M3 implements a comprehensive security framework specifically designed to address the unique challenges of AI-driven database access in medical research environments. The security architecture encompasses three critical layers of protection.
The authentication and authorization layer leverages OAuth 2.0 with JWT tokens, enabling seamless integration with standard identity providers while maintaining strict access controls. All database tools are protected by robust access control mechanisms that validate tokens according to industry best practices, ensuring that only authorized users can interact with sensitive medical data.
Query validation forms the second layer of defense through a defensive validation system that ensures only safe, read-only queries reach the database. The validator employs sophisticated syntactic analysis to automatically block potentially harmful operations, including data modification or deletion attempts, while preserving the ability to execute legitimate analytical queries essential for medical research.
The final layer implements comprehensive resource controls to maintain system stability and performance. Output limiting mechanisms limit result set sizes to prevent memory exhaustion, while rate limiting controls manage concurrent user access, ensuring consistent system responsiveness even under heavy research workloads. Together, these safeguards create a secure environment that balances accessibility with the stringent security requirements of medical data handling.
4 Results
To assess the capabilities of the M3 system, we performed an evaluation using the challenging EHRSQL 2024 benchmark [choi2024ehrsql]. This benchmark is a prominent and specialized challenge for assessing the performance of text-to-SQL systems on clinical data, using the MIMIC-IV demo database [johnson2023mimiciv_demo]. Our goal was to measure the system’s accuracy in a realistic setting and to understand the qualitative nature of its successes and failures.
4.1 Evaluation Methodology
Our evaluation dataset was derived from the official EHRSQL 2024 test set. We focused our analysis on the subset of questions deemed answerable by the dataset’s ‘is_answerable‘ flag, from which we randomly sampled 100 questions. This approach allowed us to specifically test the SQL generation and data retrieval accuracy of the system on queries where a correct answer is known to exist.
The experimental setup consisted of M3 powered by two different language models: Claude Sonnet 4 (a proprietary frontier model) [anthropic2025claude4] and gpt-oss-20B [openai_gpt_oss_20b] (an open-source model). We utilized the ‘mimic_iv.sqlite‘ database from the EHRSQL 2024 benchmark repository [ehrsql_github], which is based on MIMIC-IV demo version 2.2 [johnson2022mimiciv]. For the open-source model evaluation, we used LM Studio [lmstudio] to host gpt-oss-20B and ran tests on a MacBook M1 Max with 32GB RAM, demonstrating feasibility of local deployment on consumer hardware. We utilized the ‘mimic_iv.sqlite‘ database, which is the official database for the EHRSQL task and is based on the MIMIC-IV demo version 2.2 [johnson2022mimiciv]. This ensures that our results are directly comparable to the context of the EHRSQL benchmark.
The official EHRSQL benchmark code defines a fixed ’current time’ of "2100-12-31 23:59:00" for evaluating temporal queries [ehrsql_github]. To align our independent M3 system with this requirement, we simulated the condition by adding a contextual instruction to the start of each relevant prompt: "Set the current time to be “2100-12-31 23:59:00” when using m3 mcp." This step was essential for faithfully replicating the benchmark’s environment and validating our results.
The evaluation process involved feeding the natural language questions to the M3 system. The generated SQL queries and the final textual answers were then manually compared against the ground truth provided in the EHRSQL dataset to determine correctness.
4.2 Quantitative Performance
Out of one hundred answerable questions, both models demonstrated high accuracy. Claude Sonnet 4 correctly generated appropriate SQL queries and provided the right answer for 94 questions (94% accuracy), whereas gpt-oss-20B achieved 93% accuracy.
The comparable performance between Claude Sonnet 4 and the smaller open-source gpt-oss-20B is notable. gpt-oss-20B can be deployed locally on consumer hardware, offering a viable alternative for researchers with strict data privacy requirements or limited cloud connectivity. A detailed breakdown of the performance is presented in Table 1.
Table 1: Evaluation Results on a 100-Sample Subset of the EHRSQL Test Set
Outcome Claude Sonnet 4 gpt-oss-20B
Correct Answers 94 93
Incorrect Answers 6 7
Total Evaluated 100 100
The reported 94% accuracy was determined through a meticulous human evaluation process. For each of the 100 questions, the final answer generated by the M3 system was manually reviewed and compared against the ground truth answer from the EHRSQL dataset. An answer was deemed correct if it was logically and semantically equivalent to the ground truth, even if the phrasing or presentation differed. For example, when the ground truth is “1" (yes), automated evaluation would incorrectly penalize a response of “Yes, the patient meets this criterion" despite being semantically equivalent. This reliance on human judgment is a necessary and standard practice for evaluating complex question-answering systems, as automated scripts can fail to capture the correctness of varied but logically sound responses. This evaluation approach is consistent with methodologies used in the development of other large-scale text-to-SQL benchmarks [pourreza2023evaluating].
4.3 Visual Examples of Complex Query Results
To complement the quantitative evaluation, we also present several illustrative examples of complex queries processed by M3 on MIMIC-IV demo [johnson2023mimiciv_demo], together with their corresponding visualized outputs (Figures  3 and  4). These were generated using the MIMIC-IV demo database via the Claude-powered M3 system. Each example includes the natural language query and the resulting visualisation, designed to reflect real-world clinical insights extractable from MIMIC-IV.
Figure 3: Query: “Show trends in systolic blood pressure for patients on vasopressors within 48 hours of ICU admission.”
Figure 4: Query: “Among sepsis patients, what’s the source‑of‑infection distribution and how do groups differ in ICU stay and mortality?”
These visual outputs highlight M3’s ability not only to correctly retrieve data but also to present it in formats that are immediately interpretable to clinicians and researchers. All examples are based on the publicly accessible MIMIC-IV demo (version 2.2) and serve to illustrate the system’s practical utility in handling real-world clinical questions.
4.4 Model Comparison and Error Analysis
Analysis of the incorrect responses reveals both shared failure modes and model-specific weaknesses. Understanding these patterns provides insight into the limitations of current natural language-to-SQL systems in clinical contexts.
Common failures across both models. Three questions challenged both Claude Sonnet 4 and gpt-oss-20B, suggesting these questions contain inherent ambiguities or make unstated assumptions.
One question asked about days elapsed since a patient’s last discharge lounge stay. The gold query calculated from entry time (intime), while both models used exit time (outtime). The phrase “days since last stay” does not explicitly specify whether the reference point is the beginning or end of the stay, making both interpretations defensible.
Another question about the most frequent microbiology tests for hemodialysis patients did not specify the temporal relationship between tests and procedure. The gold query required tests to occur after hemodialysis during the same admission, but this ordering was not explicit in the question wording. Both models returned different test lists, with some overlap. The discrepancy appears to stem from how each model handled cases where multiple tests had identical frequencies, leading to different selections among equally-ranked options.
A third question asked about patients receiving serology/blood microbiology tests since 2100. Both models overcounted substantially (Claude: 41 patients, gpt-oss-20B: 40 patients) versus the gold standard (8 patients). The gold query searched for the exact specimen type ‘serology/blood’, while both models interpreted this phrase more broadly to include various blood-related tests. Without explicit enumeration of valid specimen type values, this broader interpretation represents a reasonable approach to an underspecified query.
Claude Sonnet 4-specific errors. Claude made three unique errors. A question asking for the difference between the last and second-to-last blood pressure measurements was incorrectly interpreted, with Claude comparing the first and second-to-last values instead, yielding +8.0 mmHg instead of the correct -9 mmHg. This appears to be a temporal ordering mistake in query construction.
A question about drugs prescribed after alcohol detoxification during the same hospital visit resulted in Claude identifying only phenobarbital-related medications while omitting other concurrent prescriptions (docusate sodium, haloperidol, latanoprost, omeprazole, sarna lotion). The model appears to have filtered for withdrawal-related drugs, though the question did not specify this restriction.
For a question asking for the count of current patients aged 60 or above, Claude returned 44 patients by counting those without a death date, while the correct interpretation required checking for ongoing hospitalization (dischtime IS NULL), which would yield 1 patient. This reflects the need for more precise terminology, distinguishing “currently admitted” from “alive” (not deceased) to avoid ambiguity in database queries.
gpt-oss-20B-specific errors. gpt-oss-20B made four unique errors. For a question about diagnoses following BMI 35.0-35.9 diagnosis, the model used ICD-9 code 278.00 (general obesity) as a proxy, explaining that BMI data was not directly available. However, the database actually contained a specific ICD code for “body mass index 35.0-35.9, adult” that the model failed to discover through schema exploration.
A question about patients receiving nutritional substance introduction after postprocedural pneumothorax resulted in gpt-oss-20B returning 0 patients while the ground truth indicated 1. This likely resulted from using hardcoded ICD codes rather than the gold query’s flexible approach of matching long title descriptions.
For a question asking for the first specimen test given to a patient since March 2100, gpt-oss-20B incorrectly identified “pt” (prothrombin time) from the labevents table, while the correct answer “mrsa screen” resided in the microbiologyevents table. This indicates confusion between laboratory test results and specimen collection procedures, which are tracked in different tables.
Finally, for a question asking for the top four most frequent lab tests, gpt-oss-20B’s results (glucose, chloride, sodium, hemoglobin) differed from the gold standard (chloride, creatinine, hematocrit, sodium). The discrepancy likely stems from different approaches to handling frequency ties, as the gold query uses DENSE_RANK which can return a different number of results than expected when multiple tests have the same frequency.
4.5 Discussion of Results
The evaluation results offer encouraging preliminary validation of the M3 architecture. The 93-94% accuracy demonstrates that language models, when provided with proper tools via MCP, can effectively query complex databases like MIMIC-IV without task-specific fine-tuning.
The near-equivalent performance of gpt-oss-20B has important implications. For researchers facing data privacy constraints, regulatory requirements, or limited connectivity, local deployment offers a practical path forward. The smaller parameter count contributes to lower computational requirements and faster inference on local hardware.
The qualitative analysis underscores that the primary challenge is semantic: correctly interpreting user intent.
The qualitative analysis underscores a primary remaining challenge that is not merely technical (i.e., generating valid SQL), but semantic: correctly interpreting user intent. Errors caused by linguistic ambiguity suggest future work should focus on ambiguity detection and resolution [min2020ambigqa]. The fact that both models encountered similar obstacles suggests these represent general challenges for natural language-to-SQL in clinical contexts [singhal2023large].
Most errors could potentially be resolved through multi-turn conversations where user and system iteratively refine query specification. For basic exploratory data analysis, M3 performs remarkably well. For complex analytical tasks involving nuanced temporal relationships, expert involvement remains valuable.
The transparency of M3’s approach, exposing generated SQL alongside natural language results, enables expert oversight. Researchers can verify that interpretations match intent and spot logical errors, maintaining analytical rigor.
In summary, our results indicate that the M3 system represents an important step towards simplifying access and understanding of complex clinical data. It demonstrates both high performance and, through its failures, illuminates the path forward for creating more robust and reliable natural language interfaces in the critical domain of medical research.
5 Ethical Considerations
The development and deployment of AI systems like M3 occur within societies characterized by profound forms of social, material, and political inequality. The healthcare and medical research domains are particularly susceptible to these inequalities, making it essential to address the ethical implications of technologies that democratize access to clinical data analysis [birhane2021algorithmic].
5.1 Benefits and Maintaining Analytical Rigor
M3 offers significant potential benefits by making clinical data more accessible to researchers, including those with limited computational resources or laboratory infrastructure for complex data analyses. This accessibility could advance medical knowledge and help address system-level inequalities by enabling broader participation in clinical research. The democratization of clinical data analysis represents an opportunity to engage diverse perspectives in medical research while maintaining appropriate safeguards.
The extensive training and experience previously required to conduct database queries traditionally enabled research scientists and clinicians to evaluate the scientific validity of their analyses, identify potential misinterpretations of statistical results, and understand the complexities of translating query results into clinical practice [singhal2023large]. Experienced researchers typically possess intimate knowledge of the datasets they work with, including understanding how specific categories were constructed, how data was collected, and the implications of these factors for specific research queries.
To ensure M3 users can maintain this level of analytical rigor, we recommend implementing comprehensive training programs that bridge the gap between technical accessibility and domain expertise. This includes providing detailed documentation about dataset construction, establishing mentorship programs pairing experienced researchers with new users, and creating educational resources that emphasize the importance of contextual understanding in clinical data analysis.
M3’s design principle of exposing underlying SQL queries alongside natural language results directly supports this goal by enabling users to understand and validate the analytical approach, fostering transparency and reproducibility in research workflows.
5.2 Promoting Equity and Addressing Bias
The past decade has witnessed extensive focus on bias in artificial intelligence systems, with AI algorithms shown to replicate and amplify existing forms of societal inequality and discrimination [birhane2021algorithmic]. These concerns are particularly acute in healthcare, where biased algorithms can perpetuate historical injustices and disproportionately affect communities already experiencing significant social and health inequalities [obermeyer2019dissecting].
M3 users should be equipped with training and tools to identify how specific populations, particularly marginalized groups, may be represented in datasets, and how to conduct analyses that account for potential biases. To support this, we recommend developing bias awareness training, implementing tools that help users understand the demographic composition of their analyses, and establishing review processes that evaluate research for potential equity implications.
The transparency provided by M3’s query exposition enables peer review and validation of analytical approaches, supporting the identification and correction of potential biases in research design and interpretation.
5.3 Security, Privacy, and Accountability
M3 implements comprehensive safeguards to address privacy and security concerns inherent in clinical data analysis. While MIMIC-IV consists of de-identified data, M3 maintains robust protections through comprehensive security measures. M3’s security framework includes OAuth2 authentication, query validation to prevent unauthorized operations, comprehensive audit logging, and rate limiting to prevent system abuse.
To ensure accountability in AI-assisted research, we propose a collaborative responsibility model where system developers maintain robust security measures and clear documentation, users employ the system appropriately with proper training, institutions establish governance frameworks, and the research community maintains quality standards through peer review.
The linguistic ambiguity challenges identified in our evaluation results (Section 4) highlight the importance of verification procedures. M3’s query transparency features enable users to validate their analyses and support reproducible research practices. We recommend that institutions establish procedures for reviewing AI-generated analyses, particularly those intended for clinical application or publication.
5.4 Implementation and Best Practices
To ensure M3 enhances rather than compromises scientific rigor, we recommend several best practices. Users should validate results through multiple approaches where possible, thoroughly document their analytical procedures including the natural language queries used, and ensure appropriate peer review of their work. M3’s transparency features, including exposed SQL queries and comprehensive logging, directly support these practices.
Training programs should emphasize the importance of critical evaluation skills and help users understand both the capabilities and limitations of AI-assisted analysis. By combining M3’s accessibility with robust educational frameworks, we can democratize clinical data analysis while maintaining the highest standards of scientific excellence.
Based on these considerations, we recommend institutions adopting M3 implement phased deployment strategies, beginning with supervised use in educational settings. Comprehensive training programs should address both technical and ethical aspects of AI-assisted clinical data analysis. Clear governance frameworks should establish policies for M3 usage, including guidelines for result interpretation and approval processes for sensitive analyses.
Regular monitoring of M3 usage patterns and outcomes can help identify areas for improvement and ensure alignment with institutional and professional standards. Engaging diverse stakeholders, including M3 users, clinical experts, and ethicists, will help ensure ongoing alignment with evolving best practices.
M3 represents a significant opportunity to democratize clinical data analysis while maintaining the rigor essential for advancing medical knowledge. Through careful attention to ethical considerations, comprehensive training, and robust governance frameworks, we can harness the benefits of this technology while preserving the integrity of medical research and promoting equitable access to clinical insights.
6 Conclusion and Future Work
6.1 Conclusion
M3 demonstrates that a secure, protocol-driven natural language interface to complex clinical databases is not only feasible but also highly practical for accelerating research workflows. By tightly integrating with the MIMIC-IV dataset, the system’s dual-backend architecture (supporting both local SQLite databases for rapid prototyping and cloud-scale BigQuery deployments for production research) provides flexibility for varied research settings. M3 empowers external LLM agents to perform nuanced, auditable SQL queries via self-describing tools through a two-tiered tool architecture that combines foundational database operations with domain-specific clinical functions, effectively bridging the gap between raw SQL capabilities and medical research workflows. The dual-backend design also serves an important educational function, allowing students and researchers to learn clinical data analysis techniques on local demo datasets before scaling to full production environments.
The resulting system lowers the technical and clinical barrier for researchers, enabling them to extract actionable insights from EHR data without requiring SQL expertise, schema-level familiarity, or deep knowledge of clinical workflows. Importantly, M3 preserves the security, privacy, and reproducibility required for sensitive medical data through a layered enforcement of query validation, OAuth2-based access control, and rate-limiting. These safeguards, aligned with OWASP recommendations and implemented through sqlparse-based validation and JWT token authentication, ensure that even as powerful language models gain access to clinical data backends, their queries remain constrained, interpretable, and safe.
At the same time, we acknowledge that M3 is only a starting point. Its current focus on MIMIC-IV, dependence on LLM quality, and narrow focus on data retrieval highlight opportunities for deeper integration with broader research and clinical workflows. Nonetheless, the successful deployment of M3 affirms that such interfaces can meaningfully reduce friction in data exploration, and we hope this work inspires continued development and community-driven extension.
6.2 Roadmap
We invite the research community to participate in the development of M3, submitting Pull Requests on the official github repo: https://github.com/rafiattrach/m3. Here are the list of priorities, as identified by M3 stakeholders, where we welcome immediate contributions:
A. Broader Dataset Coverage.
One of our immediate priorities is expanding M3 beyond MIMIC-IV. Planned connectors include additional PhysioNet datasets (e.g., MIMIC-CXR, MIMIC-IV-ED), multi-institutional tabular repositories like eICU, and FHIR-compatible formats. This will require a modular ingestion layer capable of abstracting over heterogeneous schemas while exposing a unified natural language interface. This expansion will be accompanied by performance optimizations including query result caching, connection pooling, and intelligent query routing to minimize latency and computational costs across diverse backend systems.
B. Richer MCP Tooling.
Future M3 versions will extend the MCP interface to include not only core SQL capabilities but also higher-level clinical tasks. These include cohort definition tools, summarization functions, declarative visualization endpoints, and retrieval-augmented generation (RAG) utilities for grounding responses in biomedical literature. Each of these will be exposed as an explicit MCP tool with well-scoped permissions.
C. Technical Enhancements.
Several technical improvements will strengthen M3’s robustness and performance. Advanced rate limiting with adaptive thresholds based on query complexity will optimize resource utilization beyond the current per-user request counting approach. Query result caching and connection pooling will improve response times for frequently accessed data patterns. Additionally, expanded authentication provider support beyond the current OAuth2/JWT implementation will accommodate diverse institutional identity management systems.
D. Ecosystem and Community Contributions.
We envision M3 evolving into a community platform for natural language–driven clinical research. To support this, we plan to introduce a plugin system and formalize contribution guidelines, including continuous integration pipelines to validate third-party ingestion, query, or analysis modules against test datasets.
Together, these enhancements will move M3 from a research prototype toward a robust, extensible foundation for secure, language-driven interaction with clinical data systems.
Acknowledgments
The authors would like to thank Dr. Gloria Hyunjung Kwak who provided clinical domain expertise, advised on cohort definitions and validation protocols. The author PM acknowledges financial support from the Fulbright Scholarship and Erasmus Mundus JM Scholarship. LAC is funded by the National Institute of Health through DS-I Africa U54 TW012043-01 and Bridge2AI OT2OD032701, the National Science Foundation through ITEST #2148451, and a grant of the Korea Health Technology R&D Project through the Korea Health Industry Development Institute (KHIDI), funded by the Ministry of Health & Welfare, Republic of Korea (grant number: RS-2024-00403047). This research was supported by a grant of the Korea Health Technology R&D Project through the Korea Health Industry Development Institute (KHIDI), funded by the Ministry of Health & Welfare, Republic of Korea (grant number: RS-2024-00439677).
M3 operates exclusively on the de‑identified MIMIC‑IV corpus released through PhysioNet under the standard Data Use Agreement. All investigators completed the required CITI “Data or Specimens Only” training, and access to the full BigQuery backend was provisioned inside credential‑bound, read‑only Google Cloud projects. No new patient‑level data were collected, stored, or exported. Lowering the technical barrier to interrogating critical‑care records carries non‑trivial misuse risks. A malicious or careless user could: (i) attempt linkage attacks that re‑identify individuals, (ii) over‑interpret associative findings as causal and deploy them for bedside decision‑making, or (iii) use the interface to generate incomplete cohorts. PhysioNet and M3 mitigate these threats through user training and with the safeguards discussed in the methodology 3.5. All empirical results, architectural diagrams, and dataset statistics cited in this manuscript were obtained from publicly available resources: the MIMIC‑IV repository on PhysioNet, the EHRSQL‑2024 benchmark materials, and the open‑source M3 codebase on GitHub.
COI Statement
The authors report no conflicts of interest.
CRediT Author Statement
Rafi Al Attrach: Investigation (lead), Software (lead), Writing – original draft (lead), Writing – review & editing (equal).
Pedro Moreira: Investigation (lead), Software (lead), Writing – original draft (lead), Funding acquisition (supporting), Writing – review & editing (equal).
Rajna Fani: Investigation (lead), Software (lead), Writing – original draft (lead), Writing – review & editing (equal).
Renato Umeton: Conceptualization (equal), Methodology (supporting), Investigation (supporting), Writing – review & editing (equal).
Amelia Fiske: Conceptualization (supporting), Writing – review & editing (equal), Ethics (lead).
Leo A. Celi: Conceptualization (equal), Supervision (lead), Project administration (lead), Writing – review & editing (equal).
This statement is based on CRediT, the ANSI/NISO Contributor Role Taxonomy.
Supporting Information
Detailed benchmark results including all natural language queries, model answers, gold standard answers, SQL queries, and error analysis notes are publicly available in the M3 GitHub repository: